OpenStudy (anonymous):

Please factor this equation. y=-2x^2-4x+3

OpenStudy (campbell_st):

well the problem that you have is that it won't factor... other than taking -1 out as a common factor

OpenStudy (raffle_snaffle):

You can expand it.

OpenStudy (anonymous):

because I had a vertex equation y-5=-2(x+1)^2 and i already changed it standard form making it y=-2x^2-4x+3 and now I want to change it to intercept form

OpenStudy (anonymous):

an example of an intercept form is (x+2)(x-3)

OpenStudy (raffle_snaffle):

nvm you can't expand it. Lol

OpenStudy (raffle_snaffle):

Ah intercept form is what I was thinking about.

OpenStudy (raffle_snaffle):

do you know how to complete the squares>

OpenStudy (anonymous):

no that was the problem

OpenStudy (raffle_snaffle):

so after you completed the squares you still were unable to change it to intercept form?

OpenStudy (anonymous):

I dont know how to complete the square

OpenStudy (raffle_snaffle):

so you take half of -4x and then square it, add it to both sides.

OpenStudy (raffle_snaffle):

http://www.youtube.com/watch?v=xGOQYTo9AKY

OpenStudy (raffle_snaffle):

Only take half of the integer. (the#)

OpenStudy (anonymous):

I got y= -2( x^2 + 2x -1.5 )

OpenStudy (raffle_snaffle):

Can you please show me your work?

OpenStudy (anonymous):

I just divided everything by -2

OpenStudy (raffle_snaffle):

Oh okay. It looks good

OpenStudy (raffle_snaffle):

Did completing the squares help out

OpenStudy (anonymous):

I dont know where to go from there

OpenStudy (math&ing001):

The formula states that (a+b)^2=a^2+2ab+b^2 In this case a=x. What you want to do is to retrieve the 'b' from the term you got (aka 2x) So 2x=2ab=2xb you get b=1 After that you add the b and you retrieve it like this: y= -2( x^2 + 2x -1.5 ) = -2( (x^2 + 2*x*1 + 1) - 1 -1.5) Can you take it from here ?

OpenStudy (anonymous):

No :(

OpenStudy (math&ing001):

y= -2( x^2 + 2x -1.5 ) = -2( (x^2 + 2*x*1 + 1) - 1 -1.5) = -2((x+1)^2 - 2.5 ) = -2(x+1)^2 +5